OpenStudy (anonymous):

You are decorating a cylindrical hat box to give to your grandmother as a gift. You want to paste family photos onto the curved part of the box. The height of the box is 8 inches and the radius is 7 inches. If none of the photos overlap, what is the area of photographs you need? Use three point one four for pi.

OpenStudy (anonymous):

2(pi)7*8=351.68 square inches

OpenStudy (anonymous):

If you unfold the cylinder, it is a rectangle with side lengths h and 2 pi r (the circumference of the circular part). A = h (2 pi r) = 2(3.14)(6)(5) A = 188.4 in^2